OpenStudy (anonymous):

PLS HELP MEDAL AND FAN PT 1 1. Is 89 prime or composite 2. Is 93 prime or composite 3. 432 is divisible by a. only 2 b. only 3 c. 2 and 3

OpenStudy (anonymous):

89 is prime. 93 is is composite. 432 is divisible by 3.

OpenStudy (anonymous):

432 is divisible by 2 and 3, sorry.

OpenStudy (anonymous):

PT 2 For questions 4-6 identify the prime factorization of the number 4. 75 a. \[25 \times 3\] b. \[2^{5} \times 3\] c. \[5^{2} \times 3\] 5. 40 a. \[2\times4\] b. \[2^{3}\] c. \[2^{2}\times5\] 6. 91 a. \[7\times13\] b. 91 x 1 c. \[3^{4}\] It's really 3 base exponent 4 times 1 but it didn't work.

OpenStudy (anonymous):

PT 3 For 7-9 identify the GCF for each given set. 7. 24 and 78 a.2 b.4 c.6 8.23 and 115 a.1 b.23 c.11 9.12 and 64 a.2 b.4 c.6

OpenStudy (anonymous):

PT 4 Identify the expression that is equivalent by the Distributive Property. 10. 5(2x + 3) a. 10 x + 3 b. 10x + 5 c. 10x + 15

OpenStudy (anonymous):

Thank you mrxbabyxo.

OpenStudy (anonymous):

\[Having trouble +stress on failing this part = me needing this answered!\]